# A Spectroscopic Search for Dormant Black Holes in Low-Metallicity Binaries

### v1

@@ Appendix A Radial Velocities
Gaia DR3 ID | JD | RV | Gaia DR3 Mean RV | Flags | RV Shift |
d | km s-1 | km s-1 | km s-1 | ||
(1) | (2) | (3) | (4) | (5) | (6) |
31530905403904512 | 2460559.9042 | ||||
31530905403904512 | 2460570.8191 | ||||
75519548132523136 | 2460569.8113 | ||||
187869776566408960 | 2460575.8128 | ||||
207547530009853952 | 2460574.8033 | ||||
249249188555702784 | 2460563.9349 | ||||
290511523477577600 | 2460553.9909 | ||||
316060462855379840 | 2460566.8837 | ||||
316671516442699520 | 2460563.9599 | ||||
327354611937489024 | 2460563.9116 | ||||
363631486266516992 | 2460568.7686 | ||||
374072654842343040 | 2460555.8470 | ||||
402635355594132224 | 2460550.9485 | ||||
462765992951416064 | 2460562.7899 | BC | |||
524686025365395328 | 2460547.9620 | hot | |||
583416851271964544 | 2460448.4926 | ||||
589980763891303040 | 2460448.5106 | ||||
596344015638431104 | 2460738.6926 | SNR | |||
614970803468896640 | 2460428.5437 | ||||
619887407151659648 | 2460450.4587 | ||||
654976435263115520 | 2460737.6410 | SNR | |||
659538102848458112 | 2460424.4863 | ||||
677442515834924416 | 2460587.0107 | ||||
677442515834924416 | 2460728.8677 | ||||
677442515834924416 | 2460760.7451 | ||||
719669191180911488 | 2460587.0275 | ||||
859588402366008320 | 2460824.7779 | ||||
866004705550987264 | 2460577.9367 | SB2 | |||
874790215772404608 | 2460577.9583 | ||||
890970938125057280 | 2460580.9244 | ||||
900812185667820160 | 2460581.9443 | ||||
979611641410430336 | 2460580.9015 | ||||
988342141692595328 | 2460577.9785 | ||||
1028448477581901184 | 2460586.0190 | ||||
1028448477581901184 | 2460717.8399 | BC | |||
1028448477581901184 | 2460728.8023 | BC | |||
1028448477581901184 | 2460760.7307 | BC | |||
1030880524287375360 | 2460585.9656 | ||||
1052266579499163392 | 2460581.9223 | ||||
1073062536268418688 | 2460581.9032 | ||||
1087217545885449600 | 2460573.8435 | BC | |||
1093810767362290560 | 2460566.9202 | ||||
1095232470257179264 | 2460575.8412 | ||||
1109010106866929536 | 2460566.9042 | ||||
1109010106866929536 | 2460582.7266 | ||||
1109010106866929536 | 2460582.7412 | ||||
1109010106866929536 | 2460584.7458 | ||||
1109010106866929536 | 2460584.7602 | ||||
1109010106866929536 | 2460584.7743 | ||||
1109010106866929536 | 2460586.7461 | ||||
1109010106866929536 | 2460586.7554 | ||||
1115628342232842368 | 2460562.7526 | ||||
1115628342232842368 | 2460704.6419 | ||||
1115628342232842368 | 2460704.6564 | ||||
1115628342232842368 | 2460704.6715 | ||||
1115628342232842368 | 2460716.8832 | ||||
1115628342232842368 | 2460716.8975 | ||||
1115628342232842368 | 2460716.9124 | ||||
1115628342232842368 | 2460716.9267 | ||||
1115628342232842368 | 2460728.8503 | ||||
1115628342232842368 | 2460735.9621 | ||||
1126103148991284352 | 2460560.7631 | ||||
1141203188852721664 | 2460562.7343 | ||||
1149756736482072576 | 2460537.7687 | ||||
1161094075752619136 | 2460424.7526 | SB2 | |||
1161094075752619136 | 2460481.6773 | SB2 | |||
1177824863156087296 | 2460479.6275 | ||||
1183864824124929792 | 2460485.6361 | ||||
1183864824124929792 | 2460738.7903 | ||||
1186550724873052160 | 2460481.5755 | ||||
1196101907506720000 | 2460480.6215 | ||||
1232281475056475392 | 2460478.5826 | ||||
1236373685536404096 | 2460515.4817 | ||||
1246628349452501120 | 2460740.7890 | SNR | |||
1255225155992171392 | 2460813.7746 | ||||
1281359310434033664 | 2460806.8174 | ||||
1283083482105602560 | 2460813.8110 | ||||
1284896817297504896 | 2460801.8186 | ||||
1290088711204610560 | 2460825.8357 | ||||
1319379563529749120 | 2460817.8264 | ||||
1324064475199832448 | 2460582.7102 | ||||
1343879087778316544 | 2460535.7197 | ||||
1360311082896731136 | 2460573.7517 | ||||
1363047492460703744 | 2460542.7441 | ||||
1373410286472677888 | 2460819.7913 | ||||
1394552589404881152 | 2460813.8255 | ||||
1397583290127619840 | 2460583.7263 | ||||
1419359598952013952 | 2460543.7390 | ||||
1422064500635586688 | 2460531.7299 | ||||
1445952348525801344 | 2460815.7786 | ||||
1475027838306979200 | 2460828.7687 | ||||
1479535389306190976 | 2460473.7443 | BC | |||
1479535389306190976 | 2460818.8051 | BC | |||
1481189192231609472 | 2460820.8115 | ||||
1489161373086670848 | 2460820.7600 | ||||
1489688863790805376 | 2460818.7585 | ||||
1497188495164830592 | 2460818.7846 | ||||
1510085903997844864 | 2460797.8060 | ||||
1512383367903955840 | 2460811.7790 | ||||
1527759075945701376 | 2460829.7525 | ||||
1542487652676277760 | 2460829.7880 | ||||
1558668134509319040 | 2460807.7762 | ||||
1558668134509319040 | 2460811.7571 | ||||
1570143840447017984 | 2460786.7921 | ||||
1574040337856735616 | 2460820.7839 | ||||
1589175287211605376 | 2460806.8404 | ||||
1598550960300560896 | 2460568.7448 | ||||
1625301043116465536 | 2460538.7154 | BC | |||
1650922176599904000 | 2460486.7830 | ||||
1650922176599904000 | 2460533.7698 | ||||
1657453619266736512 | 2460804.8068 | ||||
1663937336255868032 | 2460786.8118 | ||||
1685539544286096384 | 2460561.7649 | ||||
1694181735744784768 | 2460543.7584 | ||||
1698379915953323008 | 2460561.7148 | ||||
1711902775103746560 | 2460543.7739 | ||||
1723726369097005312 | 2460529.7376 | ||||
1731732737892580352 | 2460464.8922 | ||||
1779469856283608064 | 2460509.7781 | ||||
1788672665525489920 | 2460622.5354 | ||||
1805460872350903168 | 2460480.8230 | ||||
1810297692736381184 | 2460460.8604 | ||||
1840445472621264896 | 2460535.7920 | ||||
1843327159458210816 | 2460525.8416 | ||||
1880593731688347264 | 2460526.8949 | ||||
1908577986244357632 | 2460548.7804 | ||||
1911515915671512064 | 2460551.8352 | ||||
1919029703259598336 | 2460555.8084 | ||||
1940380874743387776 | 2460544.8901 | ||||
1974994776399961216 | 2460541.7816 | ||||
1976290172899067264 | 2460526.8769 | ||||
1985939452581432320 | 2460541.8157 | ||||
2012367039189084928 | 2460498.7167 | ||||
2012367039189084928 | 2460541.0222 | ||||
2012367039189084928 | 2460541.7998 | ||||
2038568298219233152 | 2460528.7766 | ||||
2046121840007606144 | 2460531.7873 | ||||
2047697989948528640 | 2460542.7809 | ||||
2054522692957346944 | 2460529.7828 | ||||
2074915468249960448 | 2460536.7596 | ||||
2075888368550754048 | 2460531.7708 | ||||
2096216586192095360 | 2460548.7354 | ||||
2107991153134598144 | 2460561.7451 | ||||
2118002687540997504 | 2460537.7302 | ||||
2119880309803249664 | 2460532.7342 | ||||
2123589958955565824 | 2460535.7718 | ||||
2123589958955565824 | 2460587.6801 | ||||
2123589958955565824 | 2460733.9132 | ||||
2123589958955565824 | 2460788.9790 | ||||
2123589958955565824 | 2460801.9420 | ||||
2123589958955565824 | 2460828.9601 | ||||
2178042207182395008 | 2460533.7984 | ||||
2230058934051637760 | 2460526.9145 | ||||
2236448775937348352 | 2460525.8244 | ||||
2256036476802548608 | 2460529.7530 | ||||
2281492851046104064 | 2460527.9022 | ||||
2292990925172085632 | 2460528.7569 | ||||
2306397751485602944 | 2460517.8670 | ||||
2306554049640181760 | 2460517.8837 | ||||
2315687662667389184 | 2460600.6788 | ||||
2332061310654863360 | 2460479.9237 | ||||
2344504075725183104 | 2460519.8773 | ||||
2379115460522368640 | 2460623.6257 | ||||
2400926678839820800 | 2460517.7812 | ||||
2402811860245032448 | 2460481.9391 | ||||
2402811860245032448 | 2460629.5739 | ||||
2402811860245032448 | 2460633.5548 | ||||
2409013453718653824 | 2460624.6115 | ||||
2410710034520071808 | 2460517.7977 | ||||
2429303635139814400 | 2460629.6413 | ||||
2438001734187492736 | 2460519.8280 | ||||
2438001734187492736 | 2460560.6855 | ||||
2438414532084246528 | 2460518.8925 | ||||
2474847090389656064 | 2460568.8389 | ||||
2474847090389656064 | 2460574.7516 | ||||
2474847090389656064 | 2460597.6945 | ||||
2482621256072699904 | 2460575.8227 | ||||
2487686225105416320 | 2460621.5824 | ||||
2492878703127223424 | 2460569.7471 | ||||
2511840193465032448 | 2460598.7108 | ||||
2524030925559608448 | 2460519.8610 | ||||
2554217295745049856 | 2460519.8444 | ||||
2566537942009510528 | 2460568.8061 | ||||
2570244979822205696 | 2460568.8228 | ||||
2589877859447847424 | 2460580.7619 | ||||
2613006872517192960 | 2460517.7564 | ||||
2613006872517192960 | 2460577.7052 | ||||
2613006872517192960 | 2460577.7682 | ||||
2618223166132715776 | 2460627.5910 | ||||
2619117515762724352 | 2460628.5762 | ||||
2619820000613684992 | 2460561.7327 | SNR | |||
2619820000613684992 | 2460628.5515 | ||||
2643383531229432320 | 2460510.8555 | ||||
2645540051488656768 | 2460624.5877 | SNR | |||
2667960605407262080 | 2460518.7833 | ||||
2669312489313259008 | 2460561.6874 | SNR | |||
2675008681099778944 | 2460510.7357 | ||||
2678479564430710656 | 2460623.5810 | ||||
2701466465622378496 | 2460481.8770 | ||||
2748452621862730240 | 2460629.6198 | ||||
2754870849189836800 | 2460560.7994 | ||||
2758595822787228928 | 2460627.5523 | ||||
2760474269683840256 | 2460560.7021 | ||||
2765690181047678336 | 2460519.7906 | ||||
2765690181047678336 | 2460568.7628 | ||||
2767395106968307712 | 2460624.6350 | ||||
2781500952095012096 | 2460560.8361 | ||||
2781500952095012096 | 2460567.7614 | ||||
2782781024148048000 | 2460579.6568 | ||||
2783612048780246016 | 2460575.7886 | SNR | |||
2788486767941706624 | 2460575.8055 | ||||
2788486767941706624 | 2460580.6963 | SNR | |||
2788486767941706624 | 2460596.7702 | ||||
2788486767941706624 | 2460597.6688 | ||||
2801479903005558528 | 2460555.8271 | ||||
2801532198527486976 | 2460550.9327 | ||||
2828393130155180800 | 2460629.5372 | ||||
2835800093315146496 | 2460527.9174 | ||||
2835800093315146496 | 2460544.8193 | ||||
2843092157510956032 | 2460544.8543 | ||||
2843092157510956032 | 2460548.7996 | ||||
2865540508457717376 | 2460544.8737 | ||||
2892062618706914560 | 2460625.7524 | ||||
2896172215213785344 | 2460622.7905 | SNR | |||
2896660398377207040 | 2460622.7726 | ||||
2899431683075290240 | 2460623.8294 | ||||
2905430824673932672 | 2460578.8818 | ||||
2912512023155224064 | 2460632.6903 | ||||
2949128990178535552 | 2460712.7325 | SNR | |||
2950822543021900288 | 2460712.7081 | ||||
2967475612996185728 | 2460632.6013 | SNR | |||
2972713033554244480 | 2460623.7808 | ||||
3008540860462483456 | 2460632.6359 | SNR | |||
3035481399307040384 | 2460428.5127 | ||||
3048420147209369600 | 2460732.6603 | ||||
3078328340945181952 | 2460428.4890 | ||||
3107577721076953216 | 2460735.6847 | SNR | |||
3114270895026030720 | 2460734.5719 | SNR | |||
3146951129379391232 | 2460735.6428 | SNR | |||
3150574398150536832 | 2460734.6247 | SNR | |||
3185179705568457344 | 2460626.7658 | SB2 | |||
3190523740098447744 | 2460572.8677 | ||||
3190523740098447744 | 2460574.8720 | ||||
3192689056450786304 | 2460572.8847 | SB2 | |||
3195269030422383488 | 2460622.7120 | ||||
3195269030422383488 | 2460633.5916 | SNR | |||
3205847878470249088 | 2460624.6800 | ||||
3207170625318358784 | 2460424.4686 | ||||
3214683828787855360 | 2460624.6987 | ||||
3239331546588730752 | 2460630.6720 | ||||
3246885466349457536 | 2460572.8575 | ||||
3252546886080448384 | 2460622.7369 | ||||
3258029325934167168 | 2460629.7018 | ||||
3266994193990116608 | 2460572.8197 | ||||
3270069940330680576 | 2460559.8769 | ||||
3270069940330680576 | 2460580.8054 | ||||
3289282909594133120 | 2460630.6899 | ||||
3326250307290068224 | 2460712.6276 | ||||
3334451190860930816 | 2460632.7085 | ||||
3383997379534709504 | 2460579.9028 | ||||
3449708077060728832 | 2460579.0137 | ||||
3460512226209856128 | 2460449.6214 | ||||
3463596081447832704 | 2460740.6636 | SNR | |||
3463596081447832704 | 2460740.6689 | SNR | |||
3463780112206037120 | 2460449.5844 | ||||
3485639846475600000 | 2460733.7985 | ||||
3524477002349225984 | 2460740.6931 | SNR | |||
3529854748080356352 | 2460509.4983 | ||||
3577444188310927744 | 2460427.6929 | ||||
3584915266743410048 | 2460459.5458 | ||||
3585663759283867008 | 2460479.5475 | ||||
3594336603925116672 | 2460739.7846 | SNR | |||
3618015033504305408 | 2460733.8281 | ||||
[…]

### v2

@@ Appendix A Radial Velocities
| Gaia DR3 ID | JD | RV | Instrument | Gaia DR3 Mean RV | Flags | RV Shift |
| d | km s-1 | km s-1 | km s-1 | |||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) |
| 31530905403904512 | 2460559.9042 | FEROS | ||||
| 31530905403904512 | 2460570.8191 | FEROS | ||||
| 75519548132523136 | 2460569.8113 | FEROS | ||||
| 187869776566408960 | 2460575.8128 | APF | ||||
| 207547530009853952 | 2460574.8033 | APF | ||||
| 249249188555702784 | 2460563.9349 | APF | ||||
| 290511523477577600 | 2460553.9909 | APF | ||||
| 316060462855379840 | 2460566.8837 | APF | ||||
| 316671516442699520 | 2460563.9599 | APF | ||||
| 327354611937489024 | 2460563.9116 | APF | ||||
| 363631486266516992 | 2460568.7686 | APF | ||||
| 374072654842343040 | 2460555.8470 | APF | ||||
| 402635355594132224 | 2460550.9485 | APF | ||||
| 462765992951416064 | 2460562.7899 | APF | BC | |||
| 524686025365395328 | 2460547.9620 | APF | hot | |||
| 583416851271964544 | 2460448.4926 | FEROS | ||||
| 589980763891303040 | 2460448.5106 | FEROS | ||||
| 596344015638431104 | 2460738.6926 | FEROS | SNR | |||
| 614970803468896640 | 2460428.5437 | FEROS | ||||
| 619887407151659648 | 2460450.4587 | FEROS | ||||
| 654976435263115520 | 2460737.6410 | FEROS | SNR | |||
| 659538102848458112 | 2460424.4863 | FEROS | ||||
| 677442515834924416 | 2460587.0107 | APF | ||||
| 677442515834924416 | 2460728.8677 | APF | ||||
| 677442515834924416 | 2460760.7451 | APF | ||||
| 719669191180911488 | 2460587.0275 | APF | ||||
| 859588402366008320 | 2460824.7779 | APF | ||||
| 866004705550987264 | 2460577.9367 | APF | SB2 | |||
| 874790215772404608 | 2460577.9583 | APF | ||||
| 890970938125057280 | 2460580.9244 | APF | ||||
| 900812185667820160 | 2460581.9443 | APF | ||||
| 979611641410430336 | 2460580.9015 | APF | ||||
| 988342141692595328 | 2460577.9785 | APF | ||||
| 1028448477581901184 | 2460586.0190 | APF | ||||
| 1028448477581901184 | 2460717.8399 | APF | BC | |||
| 1028448477581901184 | 2460728.8023 | APF | BC | |||
| 1028448477581901184 | 2460760.7307 | APF | BC | |||
| 1030880524287375360 | 2460585.9656 | APF | ||||
| 1052266579499163392 | 2460581.9223 | APF | ||||
| 1073062536268418688 | 2460581.9032 | APF | ||||
| 1087217545885449600 | 2460573.8435 | APF | BC | |||
| 1093810767362290560 | 2460566.9202 | APF | ||||
| 1095232470257179264 | 2460575.8412 | APF | ||||
| 1109010106866929536 | 2460566.9042 | APF | ||||
| 1109010106866929536 | 2460582.7266 | APF | ||||
| 1109010106866929536 | 2460582.7412 | APF | ||||
| 1109010106866929536 | 2460584.7458 | APF | ||||
| 1109010106866929536 | 2460584.7602 | APF | ||||
| 1109010106866929536 | 2460584.7743 | APF | ||||
| 1109010106866929536 | 2460586.7461 | APF | ||||
| 1109010106866929536 | 2460586.7554 | APF | ||||
| 1115628342232842368 | 2460562.7526 | APF | ||||
| 1115628342232842368 | 2460704.6419 | APF | ||||
| 1115628342232842368 | 2460704.6564 | APF | ||||
| 1115628342232842368 | 2460704.6715 | APF | ||||
| 1115628342232842368 | 2460716.8832 | APF | ||||
| 1115628342232842368 | 2460716.8975 | APF | ||||
| 1115628342232842368 | 2460716.9124 | APF | ||||
| 1115628342232842368 | 2460716.9267 | APF | ||||
| 1115628342232842368 | 2460728.8503 | APF | ||||
| 1115628342232842368 | 2460735.9621 | APF | ||||
| 1126103148991284352 | 2460560.7631 | APF | ||||
| 1141203188852721664 | 2460562.7343 | APF | ||||
| 1149756736482072576 | 2460537.7687 | APF | ||||
| 1161094075752619136 | 2460424.7526 | FEROS | SB2 | |||
| 1161094075752619136 | 2460481.6773 | FEROS | SB2 | |||
| 1177824863156087296 | 2460479.6275 | FEROS | ||||
| 1183864824124929792 | 2460485.6361 | FEROS | ||||
| 1183864824124929792 | 2460738.7903 | FEROS | ||||
| 1186550724873052160 | 2460481.5755 | FEROS | ||||
| 1196101907506720000 | 2460480.6215 | FEROS | ||||
| 1232281475056475392 | 2460478.5826 | FEROS | ||||
| 1236373685536404096 | 2460515.4817 | FEROS | ||||
| 1246628349452501120 | 2460740.7890 | FEROS | SNR | |||
| 1255225155992171392 | 2460813.7746 | APF | ||||
| 1281359310434033664 | 2460806.8174 | APF | ||||
| 1283083482105602560 | 2460813.8110 | APF | ||||
| 1284896817297504896 | 2460801.8186 | APF | ||||
| 1290088711204610560 | 2460825.8357 | APF | ||||
| 1294817813795066880 | 2460846.7540 | APF | ||||
| 1319379563529749120 | 2460817.8264 | APF | ||||
| 1324064475199832448 | 2460582.7102 | APF | ||||
| 1343879087778316544 | 2460535.7197 | APF | ||||
| 1360311082896731136 | 2460573.7517 | APF | ||||
| 1363047492460703744 | 2460542.7441 | APF | ||||
| 1373410286472677888 | 2460819.7913 | APF | ||||
| 1394552589404881152 | 2460813.8255 | APF | ||||
| 1397583290127619840 | 2460583.7263 | APF | ||||
| 1419359598952013952 | 2460543.7390 | APF | ||||
| 1422064500635586688 | 2460531.7299 | APF | ||||
| 1445952348525801344 | 2460815.7786 | APF | ||||
| 1475027838306979200 | 2460828.7687 | APF | ||||
| 1479535389306190976 | 2460473.7443 | APF | BC | |||
| 1479535389306190976 | 2460818.8051 | APF | BC | |||
| 1481189192231609472 | 2460820.8115 | APF | ||||
| 1489161373086670848 | 2460820.7600 | APF | ||||
| 1489688863790805376 | 2460818.7585 | APF | ||||
| 1497188495164830592 | 2460818.7846 | APF | ||||
| 1510085903997844864 | 2460797.8060 | APF | ||||
| 1512383367903955840 | 2460811.7790 | APF | ||||
| 1527759075945701376 | 2460829.7525 | APF | ||||
| 1542487652676277760 | 2460829.7880 | APF | ||||
| 1546566669018411264 | 2460862.7825 | APF | ||||
| 1558668134509319040 | 2460807.7762 | APF | ||||
| 1558668134509319040 | 2460811.7571 | APF | ||||
| 1570143840447017984 | 2460786.7921 | APF | ||||
| 1574040337856735616 | 2460820.7839 | APF | ||||
| 1589175287211605376 | 2460806.8404 | APF | ||||
| 1598550960300560896 | 2460568.7448 | APF | ||||
| 1625301043116465536 | 2460538.7154 | APF | BC | |||
| 1650922176599904000 | 2460486.7830 | APF | ||||
| 1650922176599904000 | 2460533.7698 | APF | ||||
| 1657453619266736512 | 2460804.8068 | APF | ||||
| 1663937336255868032 | 2460786.8118 | APF | ||||
| 1685539544286096384 | 2460561.7649 | APF | ||||
| 1694181735744784768 | 2460543.7584 | APF | ||||
| 1698379915953323008 | 2460561.7148 | APF | ||||
| 1711902775103746560 | 2460543.7739 | APF | ||||
| 1723726369097005312 | 2460529.7376 | APF | ||||
| 1731732737892580352 | 2460464.8922 | FEROS | ||||
| 1779469856283608064 | 2460509.7781 | FEROS | ||||
| 1788672665525489920 | 2460622.5354 | FEROS | ||||
| 1805460872350903168 | 2460480.8230 | FEROS | ||||
| 1810297692736381184 | 2460460.8604 | FEROS | ||||
| 1840445472621264896 | 2460535.7920 | APF | ||||
| 1843327159458210816 | 2460525.8416 | APF | ||||
| 1880593731688347264 | 2460526.8949 | APF | ||||
| 1908577986244357632 | 2460548.7804 | APF | ||||
| 1911515915671512064 | 2460551.8352 | APF | ||||
| 1919029703259598336 | 2460555.8084 | APF | ||||
| 1940380874743387776 | 2460544.8901 | APF | ||||
| 1974994776399961216 | 2460541.7816 | APF | ||||
| 1976290172899067264 | 2460526.8769 | APF | ||||
| 1985939452581432320 | 2460541.8157 | APF | ||||
| 2012367039189084928 | 2460498.7167 | APF | ||||
| 2012367039189084928 | 2460541.0222 | APF | ||||
| 2012367039189084928 | 2460541.7998 | APF | ||||
| 2038568298219233152 | 2460528.7766 | APF | ||||
| 2046121840007606144 | 2460531.7873 | APF | ||||
| 2047697989948528640 | 2460542.7809 | APF | ||||
| 2054522692957346944 | 2460529.7828 | APF | ||||
| 2074915468249960448 | 2460536.7596 | APF | ||||
| 2075888368550754048 | 2460531.7708 | APF | ||||
| 2096216586192095360 | 2460548.7354 | APF | ||||
| 2107991153134598144 | 2460561.7451 | APF | ||||
| 2118002687540997504 | 2460537.7302 | APF | ||||
| 2119880309803249664 | 2460532.7342 | APF | ||||
| 2123589958955565824 | 2460535.7718 | APF | ||||
| 2123589958955565824 | 2460587.6801 | APF | ||||
| 2123589958955565824 | 2460733.9132 | APF | ||||
| 2123589958955565824 | 2460788.9790 | APF | ||||
| 2123589958955565824 | 2460801.9420 | APF | ||||
| 2123589958955565824 | 2460828.9601 | APF | ||||
| 2123589958955565824 | 2460843.8971 | APF | ||||
| 2123589958955565824 | 2460854.7649 | APF | ||||
| 2123589958955565824 | 2460865.8171 | APF | ||||
| 2123589958955565824 | 2460876.8165 | APF | ||||
| 2123589958955565824 | 2460886.8202 | APF | ||||
| 2178042207182395008 | 2460533.7984 | APF | ||||
| 2230058934051637760 | 2460526.9145 | APF | ||||
| 2236448775937348352 | 2460525.8244 | APF | ||||
| 2256036476802548608 | 2460529.7530 | APF | ||||
| 2281492851046104064 | 2460527.9022 | APF | ||||
| 2292990925172085632 | 2460528.7569 | APF | ||||
| 2306397751485602944 | 2460517.8670 | FEROS | ||||
| 2306554049640181760 | 2460517.8837 | FEROS | ||||
| 2315687662667389184 | 2460600.6788 | FEROS | ||||
| 2332061310654863360 | 2460479.9237 | FEROS | ||||
| 2344504075725183104 | 2460519.8773 | FEROS | ||||
| 2379115460522368640 | 2460623.6257 | FEROS | ||||
| 2400926678839820800 | 2460517.7812 | FEROS | ||||
| 2402811860245032448 | 2460481.9391 | FEROS | ||||
| 2402811860245032448 | 2460629.5739 | FEROS | ||||
| 2402811860245032448 | 2460633.5548 | FEROS | ||||
| 2409013453718653824 | 2460624.6115 | FEROS | ||||
| 2410710034520071808 | 2460517.7977 | FEROS | ||||
| 2429303635139814400 | 2460629.6413 | FEROS | ||||
| 2438001734187492736 | 2460519.8280 | FEROS | ||||
| 2438001734187492736 | 2460560.6855 | FEROS | ||||
| 2438414532084246528 | 2460518.8925 | FEROS | ||||
| 2474847090389656064 | 2460568.8389 | FEROS | ||||
| 2474847090389656064 | 2460574.7516 | FEROS | ||||
| 2474847090389656064 | 2460597.6945 | FEROS | ||||
| 2482621256072699904 | 2460575.8227 | FEROS | ||||
| 2487686225105416320 | 2460621.5824 | FEROS | ||||
| 2492878703127223424 | 2460569.7471 | FEROS | ||||
| 2511840193465032448 | 2460598.7108 | FEROS | ||||
| 2524030925559608448 | 2460519.8610 | FEROS | ||||
| 2554217295745049856 | 2460519.8444 | FEROS | ||||
| 2566537942009510528 | 2460568.8061 | FEROS | ||||
| 2570244979822205696 | 2460568.8228 | FEROS | ||||
| 2589877859447847424 | 2460580.7619 | FEROS | ||||
| 2613006872517192960 | 2460517.7564 | FEROS | ||||
| 2613006872517192960 | 2460577.7052 | FEROS | ||||
| 2613006872517192960 | 2460577.7682 | FEROS | ||||
| 2618223166132715776 | 2460627.5910 | FEROS | ||||
| 2619117515762724352 | 2460628.5762 | FEROS | ||||
| 2619820000613684992 | 2460561.7327 | FEROS | SNR | |||
| 2619820000613684992 | 2460628.5515 | FEROS | ||||
| 2643383531229432320 | 2460510.8555 | FEROS | ||||
| 2645540051488656768 | 2460624.5877 | FEROS | SNR | |||
| 2667960605407262080 | 2460518.7833 | FEROS | ||||
| 2669312489313259008 | 2460561.6874 | FEROS | SNR | |||
| 2675008681099778944 | 2460510.7357 | FEROS | ||||
| 2678479564430710656 | 2460623.5810 | FEROS | ||||
| 2701466465622378496 | 2460481.8770 | FEROS | ||||
| 2748452621862730240 | 2460629.6198 | FEROS | ||||
| 2754870849189836800 | 2460560.7994 | FEROS | ||||
| 2758595822787228928 | 2460627.5523 | FEROS | ||||
| 2760474269683840256 | 2460560.7021 | FEROS | ||||
| 2765690181047678336 | 2460519.7906 | FEROS | ||||
| 2765690181047678336 | 2460568.7628 | FEROS | ||||
| 2767395106968307712 | 2460624.6350 | FEROS | ||||
| 2781500952095012096 | 2460560.8361 | FEROS | ||||
| 2781500952095012096 | 2460567.7614 | FEROS | ||||
| 2782781024148048000 | 2460579.6568 | FEROS | ||||
| 2783612048780246016 | 2460575.7886 | FEROS | SNR | |||
| 2788486767941706624 | 2460575.8055 | FEROS | ||||
| 2788486767941706624 | 2460580.6963 | FEROS | SNR | |||
| 2788486767941706624 | 2460596.7702 | FEROS | ||||
| 2788486767941706624 | 2460597.6688 | FEROS | ||||
| 2801479903005558528 | 2460555.8271 | APF | ||||
| 2801532198527486976 | 2460550.9327 | APF | ||||
| 2828393130155180800 | 2460629.5372 | FEROS | ||||
| 2835800093315146496 | 2460527.9174 | APF | ||||
| 2835800093315146496 | 2460544.8193 | APF | ||||
| 2843092157510956032 | 2460544.8543 | APF | ||||
| 2843092157510956032 | 2460548.7996 | APF | ||||
| 2865540508457717376 | 2460544.8737 | APF | ||||
| 2892062618706914560 | 2460625.7524 | FEROS | ||||
| 2896172215213785344 | 2460622.7905 | FEROS | SNR | |||
| 2896660398377207040 | 2460622.7726 | FEROS | ||||
| 2899431683075290240 | 2460623.8294 | FEROS | ||||
| 2905430824673932672 | 2460578.8818 | FEROS | ||||
| 2912512023155224064 | 2460632.6903 | FEROS | ||||
| 2949128990178535552 | 2460712.7325 | FEROS | SNR | |||
| 2950822543021900288 | 2460712.7081 | FEROS | ||||
| 2967475612996185728 | 2460632.6013 | FEROS | SNR | |||
| 2972713033554244480 | 2460623.7808 | FEROS | ||||
| 3008540860462483456 | 2460632.6359 | FEROS | SNR | |||
| 3035481399307040384 | 2460428.5127 | FEROS | ||||
| 3048420147209369600 | 2460732.6603 | FEROS | ||||
| 3078328340945181952 | 2460428.4890 | FEROS | ||||
| 3107577721076953216 | 2460735.6847 | FEROS | SNR | |||
| 3114270895026030720 | 2460734.5719 | FEROS | SNR | |||
| 3146951129379391232 | 2460735.6428 | FEROS | SNR | |||
| 3150574398150536832 | 2460734.6247 | FEROS | SNR | |||
| 3185179705568457344 | 2460626.7658 | FEROS | SB2 | |||
| 3190523740098447744 | 2460572.8677 | FEROS | ||||
| 3190523740098447744 | 2460574.8720 | FEROS | ||||
| 3192689056450786304 | 2460572.8847 | FEROS | SB2 | |||
| 3195269030422383488 | 2460622.7120 | FEROS | ||||
| 3195269030422383488 | 2460633.5916 | FEROS | SNR | |||
| 3205847878470249088 | 2460624.6800 | FEROS | ||||
| 3207170625318358784 | 2460424.4686 | FEROS | ||||
| 3214683828787855360 | 2460624.6987 | FEROS | ||||
| 3239331546588730752 | 2460630.6720 | FEROS | ||||
| 3246885466349457536 | 2460572.8575 | FEROS | ||||
| 3252546886080448384 | 2460622.7369 | FEROS | ||||
| 3258029325934167168 | 2460629.7018 | FEROS | ||||
| 3266994193990116608 | 2460572.8197 | FEROS | ||||
| 3270069940330680576 | 2460559.8769 | FEROS | ||||
| 3270069940330680576 | 2460580.8054 | FEROS | ||||
| 3289282909594133120 | 2460630.6899 | FEROS | ||||
| 3326250307290068224 | 2460712.6276 | FEROS | ||||
| 3334451190860930816 | 2460632.7085 | FEROS | ||||
| 3383997379534709504 | 2460579.9028 | APF | ||||
| 3449708077060728832 | 2460579.0137 | APF | ||||
| 3460512226209856128 | 2460449.6214 | FEROS | ||||
| 3463596081447832704 | 2460740.6636 | FEROS | SNR | |||
| 3463596081447832704 | 2460740.6689 | FEROS | SNR | |||
| 3463780112206037120 | 2460449.5844 | FEROS | ||||
| 3485639846475600000 | 2460733.7985 | FEROS | ||||
| 3524477002349225984 | 2460740.6931 | FEROS | SNR | |||
| 3529854748080356352 | 2460509.4983 | FEROS | ||||
| 3577444188310927744 | 2460427.6929 | FEROS | ||||
| 3584915266743410048 | 2460459.5458 | FEROS | ||||
| 3585663759283867008 | 2460479.5475 | FEROS | ||||
| 3594336603925116672 | 2460739.7846 | FEROS | SNR | |||
| 3618015033504305408 | 2460733.8281 | FEROS | ||||
| 3618107353825818880 | 2460425.7050 | FEROS | ||||
| 3618107353825818880 | 2460509.5217 | FEROS | ||||
| 3619929278952732928 | 2460481.6360 | FEROS | ||||
| 3636369696502927104 | 2460739.8331 | FEROS | ||||
| 3638399772924914176 | 2460480.5636 | FEROS | ||||
| 3651926239847699968 | 2460425.5988 | FEROS | ||||
| 3651926239847699968 | 2460509.5888 | FEROS | ||||
| 3655969212822303616 | 2460428.7253 | FEROS | ||||
| 3657930157450714752 | 2460428.6729 | FEROS | ||||
| 3679906268154294016 | 2460480.5422 | FEROS | ||||
| 3690660213428217344 | 2460462.5791 | FEROS | ||||
| 3718625668909565952 | 2460427.6692 | FEROS | ||||
| 3720871700582825088 | 2460485.6118 | FEROS | ||||
| 3736156767635254272 | 2460480.5043 | FEROS | ||||
| 3737235663419837568 | 2460462.5552 | FEROS | ||||
| 3737235663419837568 | 2460480.4853 | FEROS | ||||
| 3741881684162720768 | 2460481.4914 | FEROS | ||||
| 3816735508068512512 | 2460738.7418 | FEROS | ||||
| 3824738543969555840 | 2460449.4811 | FEROS | ||||
| 3832884821324406400 | 2460425.5153 | FEROS | SB2 | |||
| 3832884821324406400 | 2460478.4713 | FEROS | SB2 | |||
| 3838404232257203072 | 2460424.5229 | FEROS | ||||
| 3842494866483989120 | 2460449.4643 | FEROS | ||||
| 3850064076328182656 | 2460449.4980 | FEROS | ||||
[…]